Solve the following system of 3 equations. Be sure to indicate which is your working equation and which is your focus variable. 6x+18y+2z=-2 and -3x-9y+3z=17 and 12z6y-4z=-38/3.
Answer by Alan3354(69443)

Solve the following system of 3 equations. Be sure to indicate which is your working equation and which is your focus variable. 6x+18y+2z=-2 and -3x-9y+3z=17 and 12z6y-4z=-38/3.
===============
The 3rd eqn doesn't look right.
12z6y-4z=-38/3.